# Extended formulations for induced tree and path polytopes of chordal graphs

### v2

@@ 2.2 Extended formulation for the induced path polytope
The induced path cone and polytope of chordal graphs both have IDP.
@@
Proof. $(\Rightarrow)$ The extreme points of $\mathcal{P}(G)$ are scalings of $\mathcal{C}_{P}(G)$’s generators which belong to (11). Theorem 2.8 tells that the set of induced paths’ extended incidence vectors generate $\mathcal{C}_{P}(G)$ moreover, they belong to (11), which concludes the proof.
@@
Proof. $(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K\cup\{w\})]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$. By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K\cup\{w\},\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$. By summing $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$ showing that it is dominated and hence not facet-defining.
$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 11 concludes. We consider $K_{w}=K\cup\{w\}$. The points are as follows:
@@
    $\xi^{\{u,w,v\}}\ \forall uw\in\delta(w)\setminus(\delta(K)\cup\delta(C(K_{w})))$ for some vertex $v\in K$ non adjacent to $u$
The remaining $|\delta(w,C(K_{w}))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K_{w})]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$. This can be done by starting from an odd cycle and then adding the missing edges for this set to span the vertices. Now, match each edge of $\delta(w,C(K_{w}))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K_{w})|=|\delta(w,C(K_{w}))|$. Let $uw$ be in $\delta(w,C(K_{w}))$ and $uv$ be its corresponding edge of $T$. In $G$, $u$ and $v$ are non-adjacent, hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight. This construction is necessary as such a path will contain two edges of $\delta(w,C(K_{w}))$, making it impossible to continue a diagonal equal to 1. We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and "?" stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof.
$H$ is similar to the incidence matrix of the non bipartite graph induced by $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K_{w})|$.
Full row rank is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the last set of points forms a submatrix of rank $|C(K_{w})|$ which are the only ones involving elements of $\delta(w,C(K_{w}))$. This implies that the matrix has rank $|V|+|E|$, therefore the inequality is facet defining. ∎
@@
The characterization of Theorem 2.12 gives no information about the number of such facet-defining inequalities. We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently. Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, let $K_{w}=K\cup\{w\}$, $K_{w}^{\prime}=K^{\prime}\cup\{w\}$. If $K\subseteq K^{\prime}$, then we have $C(K_{w}^{\prime})\subseteq C(K_{w})$, for this reason, we say that a clique defines an orbit of $w$ if there exists no $K^{\prime}$ strictly containing $K$ such that $K_{w}\cup C(K_{w})\subseteq K_{w}^{\prime}\cup C(K_{w}^{\prime})$. Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$, inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
@@
Given a chordal graph $G$, the set of non-dominated inequalities (5) has polynomial size.
Proof. We build a chordal auxiliary graph $H$ to $G[N(w)]$, whose maximal cliques are in bijection with orbit-defining cliques of $G[N(w)]$. Polynomial size of $H$ and linear number of maximal cliques of chordal graphs will conclude.
The auxiliary graph $H$ is built from $G[N(w)]$ by first adding a true twin to each vertex of $N(w)$, let $W$ be the set of added twins. Secondly, add all the edges between elements of $W$. A perfect elimination ordering of $H$ is given by any perfect elimination of $G[N(w)]$ and followed by any permutation of the vertices of $W$. This proves that $H$ is chordal.
Let $K_{H}$ be a maximal clique of $H$. It is decomposed into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$. We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$. By construction, $K$ is a clique of $G[N(w)]$, and we show that it defines an orbit. By contradiction, suppose that a clique $K^{\prime}$ of $G[N(w)]$ strictly contains $K$ and is such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Let $K_{W^{\prime}}\subseteq W$ be the set of twins of vertices in $K^{\prime}\cup C(K^{\prime})$. By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction. Now, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows. Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction. We show that $K_{H}$ is maximal, let $v$ be a vertex of $W\setminus W_{K}$, $v$ is complete to $W_{K}$ by construction, and suppose it to also be complete to $K$. Then, its twin $u$ also is complete to $K$, implying that $u\in C(K)$ and $v\in W_{K}$, a contradiction. Let now $v\in N_{G}(w)$ be a vertex complete to $K_{H}$, then, in particular, it is complete to $W_{K}$, implying that it is complete to $C(K)$. By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit. ∎
Since for each vertex $w$, orbit-defining cliques of $G[N(w)]$ are in one-to-one correspondence with the maximal cliques of a chordal graph with at most $2|V|$ vertices, their number is $O(V)$ for fixed $w$. Summing over all $w\in V$, the total number of non-dominated inequalities (5) is $O(|V|^{2})$.
@@
There exists an extended formulation for the induced path polytope of chordal graphs with size $O(|V|^{2})$. The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
It turns out that the target of Theorems 2.12 and 2.13 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5. Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of size 4 which is bipartite therefore, Theorem 2.12 implies that its associated inequalities are not facet defining.
@@ Conclusion
In this paper, we show that the induced tree and path extended incidence vectors of chordal graphs form a Hilbert basis. We use this fact to give a compact linear system of inequalities whose binaricity characterizes chordality. It turns out that there exist non-chordal graphs for which induced tree and path extended incidence vectors form a Hilbert basis, which raises the question of characterizing such graphs. In general, showing that a set of vectors $\mathcal{V}$ forms a Hilbert basis is used to show that a linear system whose rows are elements of $\mathcal{V}$ is TDI. In this paper, we are not considering the programs whose rows are extended induced path or induced tree of chordal graphs. Finding some problems that are naturally formulated as integer linear programs whose linear constraint set has such a structure would lead to TDI-ness results.

### v3

@@ 2.2 Induced paths of chordal graphs and integral systems
The induced path cone and polytope of chordal graphs have IDP.
@@
Proof. $(\Rightarrow)$ The extreme points of $\mathcal{P}(G)$ are scalings of $\mathcal{C}_{P}(G)$’s generators which belong to (11). Theorem 2.11 tells that the set of induced paths’ extended incidence vectors generate $\mathcal{C}_{P}(G)$ moreover, they belong to (11). Finally, $\mathcal{P}(G)$ is included in $\mathcal{T}(G)$ which is bounded by Theorem 1.3. Therefore, $\mathcal{P}(G)$ is also bounded.
@@
Proof. $(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K\cup\{w\})]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$ for some $l\in\{1,\dots k\}$. By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K\cup\{w\},\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$. By summing the valid inequalities $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$. This shows that it is dominated and hence not facet-defining.
$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality for a given $w$ and $K$. This is sufficient as by equality (11) the dimension of $\mathcal{P}(G)$ is at most |V|+|E|1|V|+|E|-1. We consider $K_{w}=K\cup\{w\}$. The points are as follows:
@@
    $\xi^{\{u,w,v\}}\ \forall uw\in\delta(w)\setminus\Biggl(\delta\Bigl(K\Bigr)\cup\delta\Bigl(C(K_{w})\Bigr)\Biggr)$ for some vertex $v\in K$ non adjacent to $u$
Points 1 to 4 satisfy inequality (5) for $w$ and $K$ with equality. Moreover, they can be ordered so that the corresponding matrix is triangular. For the remaining $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ elements, two edges in $\delta\Bigl(C(K_{w})\Bigr)$ are needed to achieve equality; therefore, it is not possible to keep the triangular structure.
The remaining $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K_{w})]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$. $T_{c}$ exists and may be chosen such that $|T_{c}|=|V(\overline{G}_{c})|$: start with an odd cycle and then add the missing edges for this set to span the vertices. Now, match each edge of $\delta(w,C(K_{w}))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K_{w})|=|\delta(w,C(K_{w}))|$. Let $uw$ be in $\delta(w,C(K_{w}))$ and $uv$ be its corresponding edge of $T$. In $G$, $u$ and $v$ are non-adjacent, hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight. We refer to Figure 5 for an illustration of the matrix whose rows are the points previously exhibited and in the same order, therefore, the lines containing the submatrix $H$ correspond to the last set of $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ points. The symbol "?" stands for the submatrices associated with incidence vectors of 3-vertex paths and whose value does not need to be explicit for the validity of the proof.
$H$ is identical to the incidence matrix of the non-bipartite graph induced by $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K_{w})|$.
Full row rank of the whole matrix is obtained by the facts that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and that $H$ is the maximal submatrix with no 0 row having elements of $\delta\Bigl(w,C(K_{w})\Bigr)$ as column set. This implies that the whole matrix has rank $|V|+|E|$, therefore the inequality is facet defining. ∎
@@
The characterization of Theorem 2.15 gives no information about the number of such facet-defining inequalities. We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently.
We introduce a necessary condition on the clique $K$ for inequality 5 to be facet-defining. After that, we show that the set of cliques verifying such a property come in a polynomial number in chordal graphs.
Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, let $K_{w}=K\cup\{w\}$, $K_{w}^{\prime}=K^{\prime}\cup\{w\}$. If $K\subseteq K^{\prime}$, then we have $C(K_{w}^{\prime})\subseteq C(K_{w})$. We say that a clique defines an orbit of $w$ if there exists no $K^{\prime}$ strictly containing $K$ such that $K_{w}\cup C(K_{w})\subseteq K_{w}^{\prime}\cup C(K_{w}^{\prime})$. Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$. Inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
@@
Given a chordal graph $G$, the set of non-dominated inequalities (5) has size $O(|V|^{2})$.
Proof. We build a chordal auxiliary graph $H$ to $G[N(w)]$, whose maximal cliques are in bijection with orbit-defining cliques of $G[N(w)]$. Polynomial size of $H$ and linear number of maximal cliques of chordal graphs will concludes. The auxiliary graph $H$ is built from $G[N(w)]$ by first adding a true twin to each vertex of $N(w)$, let $W$ be the set of added twins. Secondly, add all the edges between elements of $W$.
A perfect elimination ordering of $H$ is given by any perfect elimination of $G[N(w)]$ and followed by any permutation of the vertices of $W$. This proves that $H$ is chordal.
Let $K_{H}$ be a maximal clique of $H$. It is decomposed into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$. We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$. By construction, $K$ is a clique of $G[N(w)]$, and we show that it defines an orbit. By contradiction, suppose that a clique $K^{\prime}$ of $G[N(w)]$ strictly contains $K$ and is such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Let $K_{W^{\prime}}\subseteq W$ be the set of twins of vertices in $K^{\prime}\cup C(K^{\prime})$. By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction. Conversely, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows. Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction. We show that $K_{H}$ is maximal, let $v$ be a vertex of $W\setminus W_{K}$, $v$ is complete to $W_{K}$ by construction, and suppose it to also be complete to $K$. Then, its twin $u$ also is complete to $K$, implying that $u\in C(K)$ and $v\in W_{K}$, a contradiction. Let now $v\in N_{G}(w)$ be a vertex complete to $K_{H}$, then, in particular, it is complete to $W_{K}$, implying that it is complete to $C(K)$. By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit. Since for each vertex $w$, orbit-defining cliques of $G[N(w)]$ are in one-to-one correspondence with the maximal cliques of a chordal graph with at most $2|V|$ vertices, their number is $O(V)$ for fixed $w$. Summing over all $w\in V$, the total number of non-dominated inequalities (5) is at most $O(|V|^{2})$. ∎
@@
There exists an extended formulation for the induced path polytope of chordal graphs with size $O(|V|^{2})$.
Corollary 2.18.
The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
The targets of Theorems 2.15 and 2.16 differ: there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 6. Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of length 4 which is bipartite therefore, Theorem 2.15 implies that besides defining an orbit, its associated inequalities are not facet defining.
@@ Conclusion
In this article, we show that the induced tree and path extended incidence vectors of chordal graphs respectively, form two Hilbert bases. We use this fact to give two compact linear systems of inequalities whose binaricity characterizes chordality. It turns out that there exist non-chordal graphs for which induced tree and path extended incidence vectors form a Hilbert basis, which raises the question of characterizing such graphs. These results imply that both corresponding linear optimization problems are polynomially solvable. In general, showing that some subsets of vectors of a ground set $\mathcal{V}$ forms Hilbert bases is used to show that the linear system whose rows are elements of $\mathcal{V}$ is TDI. In this work, we do not consider such systems and observe that extended incidence vectors of induced trees and path form Hilbert bases in chordal graphs. It is still a question whether these new Hilbert basis can be of any use to find new TDI systems.